OpenStudy (anonymous):

lim h->0 (ln(x+h)-ln(x))/h

OpenStudy (solomonzelman):

what is the f(x) ?

OpenStudy (solomonzelman):

you are trying to find the derivative of ln(x) using the first principles ?

OpenStudy (solomonzelman):

\(\Large\color{black}{\displaystyle\lim_{h \rightarrow ~0}\frac{\ln(x+h)-\ln(x)}{h}}\)

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

are you required to use this rule specifically, or have you learned the basic derivative such as derivative of e^x?

OpenStudy (anonymous):

I've learned "a special limit"

OpenStudy (solomonzelman):

okay, sorry for late reply. I would just maybe do it as it is, if you have to use limits. first apply, the ln(a)-ln(b) = ln(a/b).

OpenStudy (anonymous):

I did learn e^x actually, I'm just really lost on everything right now and I forgot what that was, but I'm supposed to be past that part of the book. Sorry. I just really don't understand any of this.

OpenStudy (solomonzelman):

So, we maybe can take the derivative of ln(x) as. y=ln(x) e^y=x differentiate both sides with respect to x, (dy/dx) e^y = 1 then divide both sides by e^y, and substitute x for e^y, since you know that e^y=x from the beginning.

OpenStudy (solomonzelman):

for some reason this seems easier, than doing a limit change twice.

OpenStudy (anonymous):

can you show me that...I'm really, really lost on everything.

OpenStudy (solomonzelman):

which way you want to go, do the limit?

OpenStudy (solomonzelman):

\(\Large\color{black}{\displaystyle\lim_{h \rightarrow ~0}\frac{\ln(x+h)-\ln(x)}{h}}\) \(\Large\color{black}{\displaystyle\lim_{h \rightarrow ~0}\frac{\ln(\frac{x+h}{x})}{h}}\) does this step look familiar?

OpenStudy (anonymous):

I don't know...shouldn't the ln((x+h)/2) become 1/abs((x+h)/2)

OpenStudy (solomonzelman):

why is it over 2?

OpenStudy (anonymous):

I meant x

OpenStudy (solomonzelman):

Idk, I actually haven't heard of such a rule.

OpenStudy (anonymous):

ok, probably not then

OpenStudy (solomonzelman):

I might not now all the rules, but it might be the correct rule: the way I was going to do it is: \(\Large\color{black}{\displaystyle\lim_{h \rightarrow ~0}\frac{\ln(\frac{x+h}{x})}{h}}\) \(\Large\color{black}{\displaystyle\lim_{h \rightarrow ~0}\frac{\ln(1+\frac{h}{x})}{h}}\) then, let u=h/x.... and on.

OpenStudy (solomonzelman):

not *know...

OpenStudy (anonymous):

where did you get u?

OpenStudy (solomonzelman):

I am doing substitution to change the limit.

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

I am a very slow helper apparently, maybe you can tag someone who can better help you with this?

OpenStudy (solomonzelman):

I mean it is up to you.

OpenStudy (anonymous):

no, I'm just really slow at understanding. It's not you.

OpenStudy (solomonzelman):

Well, it is a hard topic.... \(\Large\color{black}{\displaystyle\lim_{h \rightarrow ~0}\frac{\ln(1+\frac{h}{x})}{h}}\) can you tell me what do we get when we substitute u for h/x ?

OpenStudy (anonymous):

(ln(1+u))/h but what do you do with h?

OpenStudy (solomonzelman):

the h becomes ux. and h->0 gives us that u->0.

OpenStudy (solomonzelman):

\(\Large\color{black}{\displaystyle\lim_{h \rightarrow ~0}\frac{\ln(1+\frac{h}{x})}{h}}\) changing the limit, we get: \(\Large\color{black}{\displaystyle\lim_{u \rightarrow ~0}\frac{\ln(1+u)}{ux}}\)

OpenStudy (solomonzelman):

then you can take 1/x outside the limit.

OpenStudy (anonymous):

I don't know what you're saying

OpenStudy (solomonzelman):

you don't know what I mean by take the 1/x out, or don't know how I changed the limit?

OpenStudy (anonymous):

both

OpenStudy (solomonzelman):

if you don't know how to change the limit, we can take the derivative of ln(x) the way I initially proposed. \(\large\color{black}{y= \ln(x) }\) \(\large\color{black}{y= \log_e(x) }\) \(\large\color{black}{e^y=x }\) and on as I did before....

OpenStudy (anonymous):

ok, yeah

OpenStudy (solomonzelman):

So, you know that the derivative of e^x is just e^x (if you want we can prove that too)

OpenStudy (anonymous):

yeah I know that

OpenStudy (solomonzelman):

okay, but the derivative (with respect to x) of e^y will have the chain rule for (dy/dx) (dy/dx is basically same as y', so which ever notation you like better that will be what we are going to use) So differentiating e^y we get (dy/dx) * e^y, and differentiating x we get just 1. So the derivative of e^y=x is: (dy/dx) e^y = 1 then we need to solve for the dy/dx, 9dividing on e^y on both sides, (dy/dx)= 1 / e^y then we know (from the beginning) that e^y=x, so substitute x instead of "e^y". you get that; (dy/dx) = ?

OpenStudy (solomonzelman):

I'll post a prove for the derivative of e^x in just a second: while you read what i wrote.

OpenStudy (solomonzelman):

\(\Large\color{black}{\displaystyle\lim_{h \rightarrow ~0}\frac{e^{x+h}-e^x}{h}}\) \(\Large\color{black}{\displaystyle\lim_{h \rightarrow ~0}\frac{e^{x}e^h-e^x}{h}}\) \(\Large\color{black}{\displaystyle\lim_{h \rightarrow ~0}\frac{e^{x}(1-e^h)}{h}}\) \(\Large\color{black}{\displaystyle\lim_{h \rightarrow ~0}\frac{1-e^h}{h} \times \displaystyle\lim_{h \rightarrow ~0}e^{x}}\) \(\Large\color{black}{\displaystyle\lim_{h \rightarrow ~0}\frac{1-e^h}{h} \times e^{x}}\) we know that: \(\Large\color{black}{\displaystyle\lim_{h \rightarrow ~0}\frac{1-e^h}{h}=1}\) (we can prove this too if you want to) \(\Large\color{black}{1 \times e^{x}}\) \(\Large\color{black}{e^{x}}\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

(dy/dx)= 1/x

OpenStudy (solomonzelman):

yes that is right, so the derivative of ln(x) is 1/x.

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

thank you

OpenStudy (solomonzelman):

you welcome!